### v1

@@ Abstract
Adversarial attacks by malicious actors on machine learning systems, such as introducing poison triggers into training datasets, pose significant risks. The challenge in resolving such an attack arises in practice when only a subset of the poisoned data can be identified. This necessitates the development of methods to remove, i.e. unlearn, poison triggers from already trained models with only a subset of the poison data available. The requirements for this task significantly deviate from privacy-focused unlearning where all of the data to be forgotten by the model is known. Previous work has shown that the undiscovered poisoned samples lead to a failure of established unlearning methods, with only one method, Selective Synaptic Dampening (SSD), showing limited success. Even full retraining, after the removal of the identified poison, cannot address this challenge as the undiscovered poison samples lead to a reintroduction of the poison trigger in the model. Our work addresses two key challenges to advance the state of the art in poison unlearning. First, we introduce a novel outlier-resistant method, based on SSD, that significantly improves model protection and unlearning performance. Second, we introduce Poison Trigger Neutralisation ([Uncaptioned image] PTN) search, a fast, parallelisable, hyperparameter search that utilises the characteristic ”unlearning versus model protection” trade-off to find suitable hyperparameters in settings where the forget set size is unknown and the retain set is contaminated. We benchmark our contributions using ResNet-9 on CIFAR10 and WideResNet-28x10 on CIFAR100 with 0.2%, 1%, and 2% of the data poisoned and discovery shares ranging from a single sample to 100%. Experimental results show that our method heals 93.72% of poison compared to SSD with 83.41% and full retraining with 40.68%. We achieve this while also lowering the average model accuracy drop caused by unlearning from 5.68% (SSD) to 1.41% (ours).
@@ 1 Introduction
We benchmark our method against full retraining and SSD using ResNet-9 on CIFAR10 and WideResNet-28x10 on CIFAR100 with 0.2%, 1%, and 2% of the data poisoned and discovery shares ranging from a single sample all the way to 100%, in 10% increments. Our method removes 93.72% of poison compared to 83.41% of SSD, and 40.68% of retraining the model. We achieve this unlearning improvement while lowering the model accuracy drop of unlearning from 5.68% (SDD) to just 1.41% (ours). Our key contributions are:
@@
    We add a new challenge to the poison unlearning benchmark with One-Shot Healing
@@ 3.1 Outlier resistant parameter importance estimation with XLF
LF (Foster et al., 2024b) uses the computationally efficient parameter importance estimation of Aljundi et al. (2018). While this method works well for densely sampled input space regions, parameters outside this region are less reliable and can result in disproportionally low importance (Aljundi et al., 2018). In the continual learning setting of Aljundi et al. (2018) this might lead to a failure to retain these samples which causes a slight dip in model accuracy. In unlearning, on the other hand, the consequences are much more severe. In SSD-based methods, we select a parameter for dampening when the relative importance of the forget set exceeds that of the retain set times $\alpha$ ($\frac{[]_{\mathcal{S}_{forget}}}{[]_{\mathcal{S}_{retain}}}>\alpha$). Wrongly assigning a near-zero importance to a parameter for the retain set in the denominator makes the importance comparison highly susceptible to outliers caused by inaccurately estimated numerator values. This can easily lead to a wrong parameter being chosen for dampening, causing damage to the model.
@@
This is motivated by the fact that squaring the $l_{2}$ norm leads to more extreme relative values for importances. We show this in a toy example using random uniformly distributed model outputs to show the effect of squaring versus non-squaring. Fig. 3(a) shows the scaled output for $l_{2}^{w}(f(x_{k};\theta))$ with $w=[1,2]$. The importance values obtained using eq. 4 with $w=[1,2]$ in Fig. 3(b) demonstrate that the squared approach of LF produces heavier tails.
@@ 3.2 Hyperparameter search for poison contaminated data with [Uncaptioned image] PTN
    $Acc_{\mathcal{S}_{tr}}(\alpha)$: The retain set $\mathcal{S}_{tr}$ is contaminated with the undiscovered samples of $\mathcal{S}_{m}$. As we unlearn the poison trigger, the accuracy on $\mathcal{S}_{m}$ samples in $\mathcal{S}_{tr}$ will fall due to these samples being redirected/healed to their actual labels. We thus get a desired drop in accuracy that cannot be differentiated from an accuracy drop caused by model degradation.
@@ 4 Experimental setup
Our experimental setup replicates Goel et al. (2024) and adds the task of unlearning the poison trigger given a single sample out of $\mathcal{S}_{m}$ which we refer to as One-Shot Healing.
@@ 5 Results and discussion
We report the results of SSD as used in Goel et al. (2024), as well as [Uncaptioned image] PTN combined with SSD, LF, and XLF on the original benchmarking tasks and one-shot healing in Table 1. [Uncaptioned image] XLF achieves a relative improvement compared to SOTA of $\uparrow$ 12.36% on poison removal with only $24.82\%$ of relative model degradation compared to SOTA. [Uncaptioned image] XLF also sets a new SOTA for one-shot healing.

### v2

@@ Abstract
Adversarial attacks by malicious actors on machine learning systems, such as introducing poison triggers into training datasets, pose significant risks. The challenge in resolving such an attack arises in practice when only a subset of the poisoned data can be identified. This necessitates the development of methods to remove, i.e. unlearn, poison triggers from already trained models with only a subset of the poison data available. The requirements for this task significantly deviate from privacy-focused unlearning where all of the data to be forgotten by the model is known. Previous work has shown that the undiscovered poisoned samples lead to a failure of established unlearning methods, with only one method, Selective Synaptic Dampening (SSD), showing limited success. Even full retraining, after the removal of the identified poison, cannot address this challenge as the undiscovered poison samples lead to a reintroduction of the poison trigger in the model. Our work addresses two key challenges to advance the state of the art in poison unlearning. First, we introduce a novel outlier-resistant method, based on SSD, that significantly improves model protection and unlearning performance. Second, we introduce Poison Trigger Neutralisation ([Uncaptioned image] PTN) search, a fast, parallelisable, hyperparameter search that utilises the characteristic ”unlearning versus model protection” trade-off to find suitable hyperparameters in settings where the forget set size is unknown and the retain set is contaminated. We benchmark our contributions using ResNet-9 on CIFAR10 and WideResNet-28x10 on CIFAR100 with 0.2%, 1%, and 2% of the data poisoned and discovery shares ranging from a single sample to 100%. Experimental results show that our method heals 93.72% of poison compared to SSD with 83.41% and full retraining with 40.68%. We achieve this while also lowering the average model accuracy drop caused by unlearning from 5.68% (SSD) to 1.41% (ours). We further show the generalisation capabilities of our method on additional poison types with a Vision Transformer and a significantly larger dataset using ILSVRC Imagenet.
@@ 1 Introduction
We benchmark our method against full retraining and SSD using ResNet-9 on CIFAR10 and WideResNet-28x10 on CIFAR100 with 0.2%, 1%, and 2% of the data poisoned and discovery shares ranging from a single sample all the way to 100%, in 10% increments. Our method removes 93.72% of poison compared to 83.41% of SSD, and 40.68% of retraining the model. We achieve this unlearning improvement while lowering the model accuracy drop of unlearning from 5.68% (SDD) to just 1.41% (ours). We further show the performance of our method on two additional poison types and datasets using a Vistion Transformer (ViT) to show generalisation and scalability.
Our key contributions are:
@@
    We add a new challenge to the poison unlearning benchmark with One-Shot Healing and extend the poison unlearning benchmark to additional poison types, datasets, and architectures (ViT)
@@ 3.1 Outlier resistant parameter importance estimation with XLF
Figure 3: Heavy tails of near zero importance values increase the risk for wrongfully determined high relative parameter importances with noisy importance estimates. This then leads to the wrongful modification of parameters that are essential for the model’s performance and thus causes model damage.
@@
(a) Square root norm of the model output
(b) Cubed norm of the model output
Figure 5: Min-max scaled distribution comparison of taking the square-root ($l_{2}^{0.5}$) and cubing ($l_{2}^{3}$) the l2 norm of the model output for importance estimation.
@@
LF (Foster et al., 2024b) uses the computationally efficient parameter importance estimation of Aljundi et al. (2018). While this method works well for densely sampled input space regions, parameters outside this region are less reliable and can result in disproportionally low importance (Aljundi et al., 2018). In the continual learning setting of Aljundi et al. (2018) this might lead to a failure to retain these samples which causes a slight dip in model accuracy. In unlearning, on the other hand, the consequences are much more severe. In SSD-based methods, we select a parameter for dampening when the relative importance of the forget set exceeds that of the retain set times $\alpha$ ($\frac{[]_{\mathcal{S}_{forget}}}{[]_{\mathcal{S}_{retain}}}>\alpha$). Wrongly assigning a near-zero importance to a parameter for the retain set in the denominator makes the importance comparison highly susceptible to outliers caused by inaccurately estimated numerator values. This can easily lead to a wrong parameter being chosen for dampening, causing damage to the model as shown in Fig. 3.
@@
This is motivated by the fact that squaring the $l_{2}$ norm leads to more extreme relative values for importances. We show this in a toy example using random uniformly distributed model outputs to show the effect of squaring versus non-squaring. Fig. 4(a) shows the scaled output for $l_{2}^{w}(f(x_{k};\theta))$ with $w=[1,2]$. The importance values obtained using eq. 4 with $w=[1,2]$ in Fig. 4(b) demonstrate that the squared approach of LF produces heavier tails. Fig. 5 further shows that choosing a value of $w<1$ creates a heavy tail as well. The same applies to values with $w$ going beyond the original implementation of $w=2$, as shown in Fig. 5 with $w=2$, further increasing the heavy tail. Therefore, we use the most balanced option with $w=1$ for our implementation of XLF but note that in special scenarios, $w$ could be used as an adjustable parameter for potential performance gain.
@@ 3.2 Hyperparameter search for poison contaminated data with [Uncaptioned image] PTN
Figure 7: Unlearning causes an accuracy drop on the contaminated retain set due to poisoned images being redirected to their real (clean) label. This makes $\mathcal{S}_{tr}$ an unreliable metric, motivating the use of $\mathcal{S}_{f}$ for Refer to caption  to detect when the poison trigger is removed, causing a drastic drop in $\mathcal{S}_{f}$ accuracy as shown in Fig. 6.
@@
    $Acc_{\mathcal{S}_{tr}}(\alpha)$: The retain set $\mathcal{S}_{tr}$ is contaminated with the undiscovered samples of $\mathcal{S}_{m}$. As we unlearn the poison trigger, the accuracy on $\mathcal{S}_{m}$ samples in $\mathcal{S}_{tr}$ will fall due to these samples being redirected/healed to their actual labels as shown in Fig. 7. We thus get a desired drop in accuracy that cannot be differentiated from an accuracy drop caused by model degradation.
@@ 4 Experimental setup
Our initial experimental setup replicates Goel et al. (2024) and adds the task of unlearning the poison trigger given a single sample out of $\mathcal{S}_{m}$ which we refer to as One-Shot Healing. We then add a different model architecture, additional poison triggers and datasets to extend the benchmark to show the generalisation of our method to unseen tasks using the same hyperparameters as in the first task.
4.1 Goel et al. (2024) benchmarks
@@
4.2 Additional benchmarks
To show the generalisation of our methods beyond the tasks of Goel et al. (2024), we set up further experiments with additional poison attacks and datasets using a different model architecture. Notably, the method parameters will not be changed but reused from the original setting to show the robustness of our method when applied to a vastly different poison unlearning task.
We perform the additional experiments using a Vision Transformer (ViT) (Dosovitskiy et al., 2021; Steiner et al., 2021) to cover this highly relevant model architecture in modern machine learning. Furthermore, the architecture difference to ResNet allows for the demonstration of method generalisation.
All experiments are performed on Imagenette (Howard, 2019) as an additional smaller datasets, as well as the ImageNet Large Scale Visual Recognition Challenge (ILSVRC) datset (Russakovsky et al., 2015) as a large dataset going far beyond the complexity of CIFAR100.
We apply BadNet to the new datasets alongside two additional poison attacks. The first addition are sine waves overlayed on the image. We use frequency $f=0.025$ and amplitude $a=10\%$ of the maximum image value for the sine wave generation (Barni et al., 2019). The second new attack is a moving backdoor trigger, that shows the performance of unlearning methods on triggers that are not consistent amongst poisoned samples. For this poison, we place the poison trigger in a randomly chosen corner of the image, necessitating generalisation capabilities from the unlearning method to remove this poison without causing excessive harm to the model. All poison attacks are shown on an example image in Fig. 12.
The amount of poisoned samples and scenarios per dataset + poison attack combination replicates Goel et al. (2024) using $S_{m}$=[100, 500, 1000] with $S_{f}$=[0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0] $\cdot S_{m}$. We therefore extend the number of experiments by $150\%$ in relation to section 4.1.
We keep the training parameters the same as Goel et al. (2024) with the only notable changes being a lower learning rate of 0.00025 (versus 0.025) along with the use of AdamW (Loshchilov and Hutter, 2017) instead of SGD. For efficiency, we fine-tune the ViT with the poisoned dataset. We did not observe performance differences in model unlearning behaviour evaluation when training from scratch versus fine-tuning. Furthermore, fine-tuning the poison onto the model represents a realistic real-life scenario. We fine-tune the pre-trained ViT [^1] https://huggingface.co/timm/vit_base_patch16_224.augreg_in21k_ft_in1k for 2 epochs on the ILSVRC Imagenet and 5 epochs on the smaller Imagenette.
@@ 5 Results and discussion
We report the results of SSD as used in Goel et al. (2024), as well as [Uncaptioned image] PTN combined with SSD, LF, and XLF on the original benchmarking tasks and one-shot healing in Table 1. [Uncaptioned image] XLF achieves a relative improvement compared to SOTA of $\uparrow$ 12.36% on poison removal with only $24.82\%$ of relative model degradation compared to SOTA. [Uncaptioned image] XLF also sets a new SOTA for one-shot healing. We further show the generalisation capabilities of our methods on additional poison attacks and datasets using ViT, providing additional benchmarks for future works.
@@ 5.3 Computational efficiency
5.4 Further experiments with ViT on additional datasets and poison attacks
We provide additional experiments going beyond the scope of Goel et al. (2024) that demonstrate the generalisation of our method on unseen model architectures, datasets, and poison types. All experiments are performed using the same parameters used for the BadNet CIFAR experiments. Due to [Uncaptioned image] SSD consistently outperforming SSD (with compute intensive hyperparameter search), we do not report SSD results for the additional experiments and focus on the newly established SOTA approaches.
Figure 12: Data poisoning attacks
As shown in Fig. 12, we focus our discussion on four attack types that represent common attack scenarios.
  • BadNet (Gu et al., 2019) provide a simple baseline attack for evaluation.
  • Sine waves (Barni et al., 2019) are a harder to detect attack that is not overlayed onto the image as is the case with BadNet but becomes part of the image, making it harder to detect and unlearn.
  • A moving trigger (Abad et al., 2022) adds additional challenges for detection and unlearning compared to the two previous ”static” attacks.
  • Changing the label of an image without altering it poisons the dataset by harming model performance.
We do not perform experiments on the label switch attack due to its misalignment with machine unlearning. Label switching inherently requires a model to memorize the individual sample with the wrong assigned label in order to predict it as such. Therefore, there is no common link between such poisoned samples that could be unlearned that would allow for correct classification after unlearning. Goel et al. (2024) shows this in label switch experiments where half of two classes are labelled as the other. No unlearning method is able to solve this problem in a setting where $<100\%$ of the mislabelled data is identified. In that setting, the problem returns to a basic privacy focused unlearning problem. As the memorization on each example is unlearned, the sample is treated like an unseen test image and the model generalisation learned from other images in the same class correctly classifies the image (given the unlearning method did not cause excessive model damage). Schoepf et al. (2024) shows that this works in settings outside of images with a case study on unlearning mislabelled samples in a tabular setting of e-commerce order delay prediction.
Results on Imagenette with ViT on all poison types as detailed in Tab. 2, 3, and 4 show [Uncaptioned image] XLF consistently achieving the best overall results both in terms of unlearning as well as minimal model damage. It is notable, that the sinusoidal attack is the hardest to unlearn for all methods. Our hypothesis is, that it is easier to unlearn clearly separable image parts, even when these are moving, than poison that it more deeply entangled with the image itself as is the case with the sinusoidal attack.
Avg. of $S_{f}$ 10%-100% of $S_{m}$ One-shot healing ($S_{f}=1$)
Method Poison healed Model damage Poison healed Model damage
[Uncaptioned image] XLF 97.28$\pm$1.45 -0.51$\pm$0.95 72.23$\pm$14.87 -0.32$\pm$0.29
[Uncaptioned image] LF 96.60$\pm$1.90 -0.91$\pm$1.85 81.25$\pm$6.20 -0.34$\pm$0.27
[Uncaptioned image] SSD 85.99$\pm$19.04 -4.72$\pm$6.19 12.64$\pm$0.58 0.00$\pm$0.00
EU 25.82$\pm$29.34 0.18$\pm$0.13 12.74$\pm$0.87 0.02$\pm$0.03
None 12.64$\pm$0.55 0.00$\pm$0.00 12.64$\pm$0.55 0.00$\pm$0.00
Table 2: Mean results for ViT on Imagenette with BadNet attack on $S_{m}=[100,500,1000]$ with $S_{f}=[0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0]\cdot S_{m}$, reporting the percentage of poison healed compared to full retraining with all poison removed and model damage compared to the initial starting model (None). ”[Uncaptioned image] Method” denotes [Uncaptioned image] PTN search plus method.
Avg. of $S_{f}$ 10%-100% of $S_{m}$ One-shot healing ($S_{f}=1$)
Method Poison healed Model damage Poison healed Model damage
[Uncaptioned image] XLF 80.78$\pm$9.33 -2.13$\pm$2.19 23.61$\pm$5.10 -0.89$\pm$0.71
[Uncaptioned image] LF 74.16$\pm$13.79 -5.98$\pm$15.61 30.35$\pm$6.68 -2.38$\pm$2.70
[Uncaptioned image] SSD 28.75$\pm$25.65 -0.23$\pm$0.44 14.33$\pm$3.20 0.00$\pm$0.00
EU 26.15$\pm$27.18 0.34$\pm$0.24 13.33$\pm$1.12 0.31$\pm$0.18
None 14.25$\pm$2.94 0.00$\pm$0.00 14.25$\pm$2.94 0.00$\pm$0.00
Table 3: Mean results for ViT on Imagenette with a sinusoidal attack on $S_{m}=[100,500,1000]$ with $S_{f}=[0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0]\cdot S_{m}$, reporting the percentage of poison healed compared to full retraining with all poison removed and model damage compared to the initial starting model (None). ”[Uncaptioned image] Method” denotes [Uncaptioned image] PTN search plus method.
Avg. of $S_{f}$ 10%-100% of $S_{m}$
Method Poison healed Model damage
[Uncaptioned image] XLF 97.67$\pm$2.00 -1.55$\pm$1.71
[Uncaptioned image] LF 97.00$\pm$3.10 -2.19$\pm$2.79
[Uncaptioned image] SSD 74.67$\pm$12.33 -12.38$\pm$8.71
EU 65.43$\pm$14.87 -0.46$\pm$0.26
None 56.33$\pm$1.64 -0.00$\pm$0.00
Table 4: Mean results for ViT on Imagenette with moving backdoor trigger on $S_{m}=[100,500,1000]$ with $S_{f}=[0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0]\cdot S_{m}$, reporting the percentage of poison healed compared to full retraining with all poison removed and model damage compared to the initial starting model (None). ”[Uncaptioned image] Method” denotes [Uncaptioned image] PTN search plus method. No one-shot healing results are reported, as no unlearning was observed when tried.
On the larger ILSVRC dataset in Tab. 5, 6, and 7 XLF also achieves the generally best performance. The results on sinusoidal noise in Tab. 6 are to be highlighted as here SSD achieves higher unlearning but at immense cost in terms of model damage, making the unlearning results unusable and XLF the practically best method.
Comparing the results on both datasets, we can see that unlearning on the smaller dataset generally leads to more poison healing. This is to be expected, as after unlearning the poison trigger, it is easier to (re)guide the sample via model generalisation to its correct label when there are fewer classes. Learning the poison triggers also proved more difficult on the larger dataset with more classes given the numerous similar classes in ILSVRC compared to the ten classes in Imagenette. We see this behaviour as relevant for evaluation. Not having all poison triggers triggering the backdoor can make unlearning harder for methods that assume every identified modified image redirects to a malicious target. It is also unrealistic in practice to have a perfectly learned trigger, given that attackers usually only have the opportunity to poison the dataset but not change the training process.
We do not report detailed results for one-shot unlearning in Tab. 7 and 4 for the moving trigger poison, as no method was able to achieve unlearning. To achieve this, unlearning methods would need to demonstrate generalisation capabilities beyond the current SOTA.
Avg. of $S_{f}$ 10%-100% of $S_{m}$ One-shot healing ($S_{f}=1$)
Method Poison healed Model damage Poison healed Model damage
[Uncaptioned image] XLF 70.83$\pm$9.94 -0.36$\pm$0.21 38.26$\pm$38.09 -0.03$\pm$0.07
[Uncaptioned image] LF 70.23$\pm$10.01 -0.36$\pm$0.20 28.64$\pm$23.35 -0.09$\pm$0.23
[Uncaptioned image] SSD 68.13$\pm$21.04 -2.19$\pm$5.15 12.41$\pm$14.04 0.00$\pm$0.01
EU 39.80$\pm$43.46 -0.01$\pm$0.22 9.70$\pm$9.91 -0.27$\pm$0.27
None 11.02$\pm$11.62 0.00$\pm$0.00 11.02$\pm$11.62 0.00$\pm$0.00
Table 5: Mean results for ViT on Imagenet-1k (1000 classes) with BadNet attack on $S_{m}=[100,500,1000]$ with $S_{f}=[0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0]\cdot S_{m}$, reporting the percentage of poison healed compared to full retraining with all poison removed and model damage compared to the initial starting model (None). ”[Uncaptioned image] Method” denotes [Uncaptioned image] PTN search plus method.
Avg. of $S_{f}$ 10%-100% of $S_{m}$ One-shot healing ($S_{f}=1$)
Method Poison healed Model damage Poison healed Model damage
[Uncaptioned image] SSD 75.74$\pm$15.60 -6.20$\pm$5.97 43.93$\pm$42.23 0.00$\pm$0.03
[Uncaptioned image] XLF 68.00$\pm$20.43 -0.71$\pm$0.39 55.86$\pm$35.12 -2.15$\pm$3.30
[Uncaptioned image] LF 67.23$\pm$19.65 -0.81$\pm$0.56 54.62$\pm$31.60 -1.51$\pm$2.34
EU 65.11$\pm$35.64 0.12$\pm$0.26 44.46$\pm$29.18 0.13$\pm$0.27
None 43.93$\pm$36.01 0.00$\pm$0.00 43.93$\pm$36.01 0.00$\pm$0.00
Table 6: Mean results for ViT on Imagenet-1k (1000 classes) with sinusoidal attack on $S_{m}=[100,500,1000]$ with $S_{f}=[0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0]\cdot S_{m}$, reporting the percentage of poison healed compared to full retraining with all poison removed and model damage compared to the initial starting model (None). ”[Uncaptioned image] Method” denotes [Uncaptioned image] PTN search plus method.
Avg. of $S_{f}$ 10%-100% of $S_{m}$
Method Poison healed Model damage
[Uncaptioned image] XLF 96.87$\pm$1.50 -0.47$\pm$0.57
[Uncaptioned image] LF 96.85$\pm$1.59 -0.49$\pm$0.54
EU 70.60$\pm$21.26 0.22$\pm$0.37
[Uncaptioned image] SSD 69.74$\pm$17.58 -4.59$\pm$5.97
None 53.81$\pm$13.12 0.00$\pm$0.00
Table 7: Mean results for ViT on Imagenet-1k (1000 classes) with moving backdoor trigger on $S_{m}=[100,500,1000]$ with $S_{f}=[0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0]\cdot S_{m}$, reporting the percentage of poison healed compared to full retraining with all poison removed and model damage compared to the initial starting model (None). ”[Uncaptioned image] Method” denotes [Uncaptioned image] PTN search plus method. No one-shot healing results are reported, as no unlearning was observed when tried.
In summary, [Uncaptioned image] XLF achieves the best performance across the additional benchmarks using the unchanged hyperparameters from the previous experiments on the Goel et al. (2024) benchmarks. This demonstrates that the method does not require extensive hyperparameter tuning, working with the same setting from a small ResNet9 on CIFAR10 all the way to ViT on ILSVRC Imagenet with different poison types.